OpenStudy (anonymous):

Directions: Type the correct answer in each box. Spell all words correctly, and use numerals instead of words for numbers. If necessary, use decimals instead of fractions. Mrs. Matthews came up with an expression. Holly and Jim wrote the expression in their notebooks in order to simplify the expression. They compared their expressions. Holly's Notebook Jim's Notebook Holly's expression is equal to , and Jim's expression is equal to . So, 's expression is less than the other expres

OpenStudy (anonymous):

I think this is different

OpenStudy (misty1212):

i'll say !!

OpenStudy (anonymous):

Holly's Notebook \[6 x (24 / 2)^2\]

OpenStudy (misty1212):

what the monkey kind of on line system is this?

OpenStudy (anonymous):

STUDY ISLAND XD

OpenStudy (anonymous):

I wanna kill them but oh well

OpenStudy (misty1212):

lolol "study island" indeed is it FLVS or sommat?

OpenStudy (anonymous):

No clue. Just math 0-0

OpenStudy (misty1212):

i mean are you going to on line school?

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

dont they have real schools where you are?

OpenStudy (anonymous):

Jim's Notebook \[6 x 24 \]

OpenStudy (misty1212):

\[6\times 24\]

OpenStudy (anonymous):

They do. But they are crappy

OpenStudy (anonymous):

WTF i put the divisor in

OpenStudy (misty1212):

as crappy as this?

OpenStudy (misty1212):

\[6\times \frac{24}{2}\]

OpenStudy (anonymous):

\[Jim's Notbook 6 x 24 / 2^2\]

OpenStudy (anonymous):

FAIL

OpenStudy (anonymous):

6 x 24 / 2^2

OpenStudy (misty1212):

lolol you are adorable \[6\times \frac{24}{2^2}\]

OpenStudy (anonymous):

Holly's notebook 6 (24 / 2)^2

OpenStudy (misty1212):

like that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You need to compare holly's and jim's notebook. XD

OpenStudy (misty1212):

ok then \[\huge 6\times \left(\frac{24}{2}\right)^2\]

OpenStudy (misty1212):

is that holly's note book ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

jim has no parthensis on his

OpenStudy (anonymous):

So his would be. 6 x 24 / 2^2

OpenStudy (anonymous):

Any clue what it is?

OpenStudy (misty1212):

yes i do

OpenStudy (anonymous):

:D

OpenStudy (misty1212):

first square 2

OpenStudy (misty1212):

you get \[6\times \frac{24}{4}\]

OpenStudy (misty1212):

then divide, to get \[6\times 6\]

OpenStudy (misty1212):

which is \(36\)

OpenStudy (anonymous):

So I put in the first square is 2. Holly's Notebook, correct?

OpenStudy (anonymous):

And true

OpenStudy (misty1212):

no that was jim's notebook

OpenStudy (misty1212):

holly's you have to divide first, then square

OpenStudy (anonymous):

ah

OpenStudy (misty1212):

\[6\left \frac{24}{2}\right)^2=6\times 12^2=6\tiems 144\]

OpenStudy (anonymous):

[]=[]

OpenStudy (misty1212):

\[6\left( \frac{24}{2}\right)^2=6\times 12^2=6\times 144\]

OpenStudy (anonymous):

OH

OpenStudy (misty1212):

that was holly's

OpenStudy (anonymous):

6 x 144 is holly's

OpenStudy (anonymous):

What is jim's?

OpenStudy (misty1212):

you can compute that if you like

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

back

OpenStudy (anonymous):

After this can you help me with an english assignment?

OpenStudy (anonymous):

Yes no?